An editor left home at 7:15 AM and took a bus to the office. The bus averaged 27 mph and arrived at the office building at 7:55 AM. At 5:00 PM, a co-worker drove the editor home, where she arrived at 5:24 PM. What was the average speed for the drive home?
Answer by htmentor(1343)

The average speed of the bus = 27 mph. The travel time to work is
55 - 15 = 40 mins.
Thus the distance to work = d = v*t = 27 mi/h * (40/60)h = 18 miles
The drive home took 24/60 = 2/5 h, so the average speed = 18/(2/5) = 45 mph